OpenStudy (anonymous):

Jeremy bought a new truck for $32,000. The value of the truck after t years can be represented by the formula V = 32,000(.8)t. When will the truck be worth approximately $6700?

OpenStudy (anonymous):

I'm assuming you mean for the equation to be \[V=32,000(.8)^t\] Because it doesn't work if it is a linear equation. So if V represents the value of the truck then with our equation \[V=32,0000(.8)^t\] We plug in 6,700 for V and solve for t, because V is the future value of the truck, and t is the number of years after purchase of the truck that it will have the value of $6,700. So, \[6700=32,000(.8)^t\] We simplify by dividing both sides by 32,000 to get: \[0.18108=(0.8)^t\] From here we take the common, base 10 logarithm of both sides (though any other base works, as long as you log both sides with the same base). \[log(0.18108)=t log(0.8)\] The t moves from a power to coefficient of log(0.8) because a logarithm "undoes" a exponential function (this is also a property of logs). From there, we divide both sides by log(0.8) to get: \[\frac {\log(0.18108)}{\log(0.8)}=t\] So then, \[t=7.66\] But years can only be whole numbers, so we round up to \[t=8\]

OpenStudy (anonymous):

Would this be an example of exponential decay?

OpenStudy (anonymous):

Yes, because as time goes on, the value of something decreases, or decays.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem :)